OpenStudy (anonymous):

I have 5 question left on my homework sheet which is do tonight. My little girl is not allowing me much time to get the work done. Can you guys please help. I have to show all work.

OpenStudy (anonymous):

shoot.....

OpenStudy (anonymous):

questions?

OpenStudy (anonymous):

Eric is five years older than his sister Lisa, while Shawn is two years younger the Lisa. Write an expression for the sum of their ages.

OpenStudy (anonymous):

3x-5/2=-1/4

OpenStudy (anonymous):

0.3(4x+5)-1.7(x-4)=6.8

OpenStudy (anonymous):

let Lisa's age = x years then Eric's age = (x + 5) years and Shawn's age = (x - 2) years sum of their ages = x + (x+5) + (x-2) = 3x + 3

OpenStudy (anonymous):

Next pls...

OpenStudy (anonymous):

L=2b{3}/a find L when a=6 and b=3

OpenStudy (anonymous):

substitute the values in the given equation 2*3{3} 18 L = ------- = ----- = 3 6 6

OpenStudy (anonymous):

next....

OpenStudy (anonymous):

v=1/3pier{2}h solve for h

OpenStudy (anonymous):

{2} --- is it "raised to power 2" ???? and was {3} in previous question "raised to power 3" ???

OpenStudy (anonymous):

Was previous question as follows ??? \[L=2b ^{3}/a\]

OpenStudy (anonymous):

and last question\[V=1/3*\pi*r ^{2}*h\]

OpenStudy (anonymous):

pls clarify...

OpenStudy (anonymous):

Sorry little one demands attewntion Yes they are raised to the 2 and 3 power as you have

OpenStudy (anonymous):

we use the carat sing^ to indicate power so x raised to power five is shown as x^5 pls use the same convention to avoid confusion n mistakes.. so I send the correct solution for the previous one....... and then for the latest one....

OpenStudy (anonymous):

L=2b^3/a find L when a=6 and b=3 2*3^3 2*27 54 L = ------ = ------ = ------ = 9 6 6 6

OpenStudy (anonymous):

v=1/3pier{2}h solve for h π * r^2 * h V = ----------- 3 3V = π * r^2 * h 3V ------- = h π * r^2

OpenStudy (anonymous):

next pls.... if u hv to show raised to power, use the sign "^" to indicate power...

OpenStudy (anonymous):

3x-5-2=-1/4

OpenStudy (anonymous):

I am lost on the question with pie

OpenStudy (anonymous):

3x - 7 = -1/4 multiply both sides by 4 (this is done to cancel out the 4 in the denominator of -1/4) 4(3x - 7) = 4(-1/4) 12x - 28 = -1 12x = -1 + 28 x = 27/12

OpenStudy (anonymous):

what hv u not understood in the previous question with π in it ????

OpenStudy (anonymous):

None of it

OpenStudy (anonymous):

do u hv a SKYPE account as this will be easier to explain if v can hv a voice chat my SKYPE id is hakki.singh

OpenStudy (anonymous):

And the other question I plued in wrong it should have been 3x-5/2=-1/4 I would first need to find the LCD right which is 4 and then muliply the left side by 2 and get 6x-10/4=-1/4

OpenStudy (anonymous):

the simpler method is as follows:..... 3x-5/2=-1/4 5 -1 3x - ------ = ------ 2 4 Here 4 is multiple of 2, so multiply both sides by 4 (means multiply all terms by 4) 5 -1 4(3x) - 4*----- = 4*----- 2 4 u get 12x - 10 = -1 (you hv got rid of both denominators and no need to find LCD !!!!) 12x = -1 + 10 x = 9/12 x = 3/4

OpenStudy (anonymous):

I do not have a skype account